OpenStudy (cosma09):

Which shape has 2 pairs of congruent sides and 4 right angles? A. rectangle B. trapezoid C. triangle D. pentagon

OpenStudy (cosma09):

@Conner_Isom-224-1091

OpenStudy (anonymous):



OpenStudy (cosma09):

hi

OpenStudy (anonymous):

you have the answer already lol

OpenStudy (cosma09):

I know

OpenStudy (cosma09):

Which equation could be used to determine the area of a rectangle that is 9 cm long and 13 cm wide? A. A = 2 × (9 × 13) B. A = 2 × (9 + 13) C. A = 9 × 13 D. A = 9 + 13